# QKFormer: Hierarchical Spiking Transformer using Q-K Attention

### v1

@@ 3.3 QKFormer
3.3.1 Overall Hierarchical Architecture
The overview of QKFormer architecture is presented in Figure 2. The input form can be formulated as $(T_{0}\times H\times W\times n)$. In static RGB image datasets, $T_{0}=1$ and $n=3$. In temporal neuromorphic datasets, the input $T_{0}=T$, while $n=2$.
In our implementation, we use a patch size of $4\times 4$ and thus the input feature dimension ($4\times 4\times n$) of each patch is projected into a spike-form arbitrary dimension (denoted as $C$) in Patch-Embed 1, which together with the following QKFormer blocks are referred to as ”Stage 1”. The number of tokens in Satge 1 is $(\frac{H}{4}\times\frac{W}{4})$.
To produce a hierarchical spiking representation, the number of tokens is reduced in Patch-Embed 2 and Patch-Embed 3 as the network goes deeper. Both Patch-Embed 2 and Patch-Embed 3 reduce the number of tokens by a patch size of $2\times 2$ (2×2$\times$2 × downsampling of resolution), and the number of channels is transformed into $2C$ and $4C$, respectively. We denote the Patch-Embed 2 and the following QKFormer blocks as ”Stage 2”, which reduces the number of tokens $(\frac{H}{8}\times\frac{W}{8})$. While Patch-Embed 3 and the following Spikformer blocks are referred to as ”Stage 3” with $(\frac{H}{16}\times\frac{W}{16})$ tokens. The number of spiking transformer blocks (QKFormer or Spikformer) in each stage are $N_{1}$, $N_{2}$, and $N_{3}$, respectively. These stages jointly produce a hierarchical spike-form representation.
Similar to the standard transformer encoder block, a QKFormer block contains a Q-K Attention module (QKTA or QKCA) and a Spiking MLP (SMLP) block, which can be formulated as follows:
$X_{l}^{\prime}=\operatorname{QKTA}\left(X_{l-1}\right)+X_{l-1},X_{l}^{\prime}\in\mathbf{R}^{T\times N\times D},$ (13)
$X_{l}=\operatorname{SMLP}\left(X_{l}^{\prime}\right)+X_{l}^{\prime},X_{l}\in\mathbf{R}^{T\times N\times D}.$ (14)
At last, a fully connected layer (FC) is used for classification.
$Y=\operatorname{FC}\left(\operatorname{GAP}\left(X_{L}\right)\right),$ (15)
where $\operatorname{GAP}$ means global average pooling.
@@ 3.3.1 Overall Hierarchical Architecture
Figure 2: The overview of QKFormer, a hierarchical spiking transformer with Q-K attention. The QKFormer block is shown on the right panel.
Table 2: Results on ImageNet-1k classification. ${}^{\ast}$ means $288^{2}$ input resolution when inference, ${}^{\ast\ast}$ means $384^{2}$ input and others are $224^{2}$ by default. The top-5 accuracy of QKFormer (${}^{\ast\ast}$) is 97.74%.
@@
ImageNet-1K SEW ResNetFang et al. (2021a) SEW-ResNet-34 21.79 4 67.04
SEW-ResNet-101 44.55 4 68.76
SEW-ResNet-152 60.19 4 69.26
MS-ResNetHu et al. (2021) MS-ResNet-18 11.69 6 63.10
MS-ResNet-34 21.80 6 69.42
${}^{\ast}$ 77.28 5 76.02
SpikformerZhou et al. (2023c) Spikformer-8-384 16.81 4 70.24
Spikformer-8-512 29.68 4 73.38
Spikformer-8-768 66.34 4 74.81
SpikingformerZhou et al. (2023a) Spikingformer-8-384 16.81 4 72.45
Spikingformer-8-512 29.68 4 74.79
Spikingformer-8-768 66.34 4 75.85
S-TransformerYao et al. (2023) S-Transformer-8-384 16.81 4 72.28
S-Transformer-8-512 29.68 1 71.68
S-Transformer-8-512 29.68 4 74.57
${}^{\ast}$ 66.34 4 77.07
CMLZhou et al. (2023b) Spikformer-8-384 16.81 4 72.73
Spikformer-8-512 29.68 4 75.61
Spikformer-8-768 66.34 4 77.34
QKFormer HST-10-384 16.47 4 78.80
HST-10-512 29.08 4 82.04
HST-10-768 64.96 1 81.69
HST-10-768 64.96 4 84.22
${}^{\ast}$ 64.96 4 85.25
${}^{\ast\ast}$ 64.96 4 85.65
3.3.2 Patch Embedding with Deformed Shortcut.
The input and output of a patch embedding block have different channel and token numbers. To realize residual learning in patch embedding, we can perform a lightweight linear projection $W_{d}$ in the shortcut connections to match the channel and token numbers . Given the input spiking map $\mathbf{X}$, the process of patch embedding can be formulated as follows:
$\mathbf{Y}=\mathcal{F}\left(\mathbf{X},\left\{W_{i}\right\}\right)+\mathrm{SN}(W_{d}\mathbf{X}).$ (16)
[…]
There are mainly two types of residual shortcuts in deep SNNs. Formula.16 shows the patch embedding in the way of activation-before-addition Fang et al. (2021a); Zhou et al. (2023c). The other way of the patch embedding with the pre-activation residual shortcut Hu et al. (2021); Zhou et al. (2023a); Yao et al. (2023) can be formulated as follows:
$\mathbf{Y}=\mathrm{SN}(\mathcal{G}\left(\mathbf{X},\left\{W_{j}\right\}\right)+W_{d}\mathbf{X}),$ (17)
@@
Table 3: Comparision on CIFAR10, CIFAR100, DVS128, CIFAR10-DVS.
@@
CIFAR10 SpikformerZhou et al. (2023c) Spikformer-4-384 9.32 4 95.51
SpikingformerZhou et al. (2023a) Spikingformer-4-384 9.32 4 95.81
CMLZhou et al. (2023b) Spikformer-4-384 9.32 4 96.04
S-TransformerYao et al. (2023) S-Transformer-2-512 10.28 4 95.60
QKFormer HST-4-384 6.74 4 96.18
CIFAR100 SpikformerZhou et al. (2023c) Spikformer-4-384 9.32 4 78.21
SpikingformerZhou et al. (2023a) Spikingformer-4-384 9.32 4 78.21
CMLZhou et al. (2023b) Spikformer-4-384 9.32 4 80.02
S-TransformerYao et al. (2023) S-Transformer-2-512 10.28 4 78.4
QKFormer HST-4-384 6.74 4 81.15
DVS128 SpikformerZhou et al. (2023c) Spikformer-2-256 2.57 10 / 16 96.9 / 98.3
SpikingformerZhou et al. (2023a) Spikingformer-2-256 2.57 10 / 16 96.2 / 98.3
CMLZhou et al. (2023b) Spikformer-2-256 2.57 10 / 16 97.6 / 98.6
S-TransformerYao et al. (2023) S-Transformer-2-256 2.57 - / 16 - / 99.3
STSAWang et al. (2023) STSFormer-2-256 1.99 10 / 16 97.3 / 98.7
QKFormer HST-2-256 1.50 10 / 16 98.3 / 98.6
CIFAR10-DVS SpikformerZhou et al. (2023c) Spikformer-2-256 2.57 10 / 16 78.9 / 80.9
SpikingformerZhou et al. (2023a) Spikingformer-2-256 2.57 10 / 16 79.9 / 81.3
CMLZhou et al. (2023b) Spikformer-2-256 2.57 10 / 16 79.2 / 80.9
S-TransformerYao et al. (2023) S-Transformer-2-256 2.57 - / 16 - / 80.0
STSAWang et al. (2023) STSFormer-2-256 1.99 10 / 16 78.96 / 79.93
QKFormer HST-2-256 1.50 10 / 16 83.8 / 84.0
@@ 4 Experiments
In this section, we first evaluate the classification performance of our proposed QKFormer on the large-scale image dataset, ImageNet-1K Deng et al. (2009). Then, we assess the performance of QKFormer on small-scale static datasets, CIFAR Krizhevsky (2009) (including CIFAR10 and CIFAR100). Furthermore, we examine QKFormer on popular temporal neuromorphic datasets, including CIFAR10-DVS Li et al. (2017) and DVS128 Gesture Amir et al. (2017). Finally, we conduct a series of ablation studies to further understand the model.
@@ 4.1 Results on ImageNet-1k Classification
In this experiment, we use AdamW as the optimizer, which is adopted with a base learning rate of $6\times 10^{-4}$. The actual learning rate was calculated as BatchSize/256 multiplied by the base learning rate. The batch size is set to $512$, which is realized by accumulated gradient iterations He et al. (2022) and distributed across 8 Nvidia V100 GPUs. We trained QKFormer for $200$ epochs. In addition, following DeiT Touvron et al. (2021), data augmentation techniques including RandAugment Cubuk et al. (2020), random erasing Zhong et al. (2020), and stochastic depth Huang et al. (2016) are employed in this study. The number of blocks in the three stages is set as {1, 2, 7} respectively. The architecture of QKFormer is named Hierarchical Spiking Transformer (HST). The ImageNet-1k classification results of QKFormer are shown in Table 2. Note that SEW ResNet, Spikformer, CML and our QKFormer in this experiment are all based on activation-before-addition shortcuts, while MS ResNet, Spikingformer, and S-Transformer are all based on the pre-activation shortcuts.
@@
The experimental results demonstrate the superior performance of our proposed QKFormer, surpassing previous works’ performance by a large margin (Table 2). QKFormer (64.96 M) achieves 85.65% top-1 accuracy and 97.74% top-5 accuracy on ImageNet. To begin with, we compare our model with the baseline spiking transformer (i.e., Spikformer Zhou et al. (2023c)). Our QKFormer models have slightly fewer parameters but much higher performance. For example, our QKFormer (64.96 M, 85.65%) significantly outperforms Spikformer (66.34 M, 74.81%) by 10.84%. In addition, compared with SDSA, our Q-K attention has lower computational complexity (shown in Table 1) and our QKFormer has much higher performance than S-Transformer (built by SDSA) Yao et al. (2023). In detail, QKFormer outperforms S-Transformer by 7.55%, 7.47%, and 8.58% respectively on three models with comparable #parameters. Finally, Our QKFormer outperforms the SOTA model CML Zhou et al. (2023b) by 6.07%, 6.43%, and 8.31% respectively on three models with comparable #parameters. To our best knowledge, this is the first time that a direct training SNN model has achieved an accuracy of over 85% on ImageNet-1k.
4.2 Results on Small Dataset Classification
We evaluate our QKFormer on small-scale datasets, including CIFAR10, CIFAR100 Krizhevsky (2009) and temporal neuromorphic datasets (CIFAR10-DVS and DVS128 Gesture Amir et al. (2017)). The overview results on the four small-scale datasets are presented in Table 3.
4.2.1 CIFAR Classification
@@
In this experiment, the QKFormer is trained for 400 epochs with a batch size of 64 following previous works: Spikformer Zhou et al. (2023c), Spikingformer Zhou et al. (2023a). Following Spikformer, we use 4 blocks in QKFormer in total, which are distributed {1, 1, 2} in three stages. Due to the hierarchical architecture design, our QKFormer model has only 6.74 M parameters in this case.
The results on CIFAR datasets are shown in Table 3. For CIFAR10, our model achieved 96.18% accuracy with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 0.67% and reduces 2.58 M parameters meanwhile. For CIFAR100, our model achieved 81.15% with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 2.94% and reduces 2.58 M parameters meanwhile.
4.2.2 Temporal Neuromorphic Classification
@@
We compare our method with SOTA methods on both CIFAR10-DVS and DVS-Gesture datasets. In this experiment, We utilize a mini QKFormer model with 1.50 M parameter, which has {0, 1, 1} blocks in three stages. The max patch embedding dimension is set to 256. The training process involves 200 epochs for DVS128 Gesture and 106 epochs for CIFAR10-DVS. The number of time steps of the spiking neuron is 10 or 16.
The experimental results of temporal neuromorphic classification are presented in Table 3. For DVS128-Gesture dataset, our model with 1.50 M parameters achieves 98.6% accuracy using 16 time steps and 98.3% accuracy using 10 time steps. For CIFAR10-DVS dataset, our model achieves 84.0% accuracy with only 1.50 M parameters using 16 time steps. Our proposed QKFormer significantly outperforms Spikformer by 3.1% while reducing 1.07 M parameters. In addition, our model with 10 time steps achieves 83.8% accuracy, which outperforms Spikformer by 4.9% and outperforms the SOTA model (Spikingformer) by 3.9%.

### v2

@@ 3.4 QKFormer
Figure 2: The overview of QKFormer, a hierarchical spiking transformer with Q-K attention.
Overall Hierarchical Architecture. The overview of QKFormer is presented in Figure 2. The input form can be formulated as $(T_{0}\times H\times W\times n)$. In static RGB image datasets, $T_{0}=1$ and $n=3$. In temporal neuromorphic datasets, the input $T_{0}=T$, while $n=2$. In our implementation, we use a patch size of $4\times 4$ and thus the input feature dimension ($4\times 4\times n$) of each patch is projected into a spike-form arbitrary dimension (denoted as $C$) in Spiking Patch Embedding with Deformed Shortcut 1 (SPEDS-1), which together with the following QKFormer blocks are referred to as "Stage 1". The number of tokens in Satge 1 is $(\frac{H}{4}\times\frac{W}{4})$. To produce a hierarchical spiking representation, the number of tokens is reduced in SPEDS-2 and SPEDS-3 as the network goes deeper. Both SPEDS-2 and SPEDS-3 reduce the number of tokens by a patch size of $2\times 2$ (2×2$\times$2 × downsampling of resolution), and the number of channels is transformed into $2C$ and $4C$, respectively. We denote the SPEDS-2 and the following QKFormer blocks as "Stage 2", which reduces the number of tokens $(\frac{H}{8}\times\frac{W}{8})$. While SPEDS-3 and the following Spikformer or QKormer blocks are referred to as "Stage 3" with $(\frac{H}{16}\times\frac{W}{16})$ tokens. The number of spiking transformer blocks (QKFormer or Spikformer) in each stage are $N_{1}$, $N_{2}$, and $N_{3}$, respectively. These stages jointly produce a hierarchical spike-form representation.
Mixed Spiking Attention Integration. In the former stage of a hierarchical architecture model, the number of channels is small while the number of tokens is large. In the latter stage, the channel number is large while the token number is small. Thus it leads to suboptimal performance when we only use a single type of Q-K attention in a hierarchical architecture model. Therefore, we use mixed spiking attention integration in QKFormer. QKTA is conducted in the former stage in hierarchical architecture, and we could choose QKCA or SSA in the latter stage. In the subsequent experiments, we use SSA in the last stage of QKFormer and QKTA in the former stages by default.
QKFormer Blocks. Similar to the standard transformer encoder block, a QKFormer block contains a Q-K Attention module (QKTA or QKCA) and a Spiking MLP (SMLP) block, which can be formulated as follows:
$\mathbf{X}_{l}^{\prime}=\operatorname{QKTA}\left(\mathbf{X}_{l-1}\right)+\mathbf{X}_{l-1},\mathbf{X}_{l}^{\prime}\in R^{T\times N\times D},$ (8)
$\mathbf{X}_{l}=\operatorname{SMLP}\left(\mathbf{X}_{l}^{\prime}\right)+\mathbf{X}_{l}^{\prime},\mathbf{X}_{l}\in R^{T\times N\times D}.$ (9)
At last, a fully connected layer is used as the classifier behind the last block.
3.5 Spiking Patch Embedding with Deformed Shortcut.
Residual shortcuts in SNNs [28] can implement identity mapping, which reduces information loss (facilitates information transmission and integration) in spike communication, thus ensuring the network can be well-behaved in a depth-insensitive way. Previous spiking transformers [11, 12, 13] use the residual shortcuts to achieve identity mapping, mainly focusing on the spiking attention block and spiking MLP block, and lacking identity mapping in patch embedding across the downsampling block. The input and output of a spiking patch embedding block in QKFormer have different channel and token numbers. To realize residual learning in spiking patch embedding, we can perform a lightweight linear projection $\mathbf{W}_{d}$ in the shortcut connections to match the channel and token numbers, thus realizing the identity mapping cross downsampling blocks in spiking patch embedding. Given the input spiking map $\mathbf{X}$, the process of patch embedding can be formulated as follows:
$\mathbf{Y}=\mathcal{F}\left(\mathbf{X},\left\{\mathbf{W}_{i}\right\}\right)+\mathrm{SN}(\mathbf{W}_{d}\mathbf{X}).$ (10)
@@
There are mainly two types of residual shortcuts in deep SNNs. Formula.10 shows the patch embedding in the way of activation-before-addition [28, 11]. The other way of the patch embedding with the pre-activation residual shortcut [31, 12, 13] can be formulated as follows:
$\mathbf{Y}=\mathrm{SN}(\mathcal{G}\left(\mathbf{X},\left\{\mathbf{W}_{j}\right\}\right)+\mathbf{W}_{d}\mathbf{X}),$ (11)
@@
Table 2: Results on ImageNet-1K. Power is calculated as the average theoretical energy consumption when predicting an image from ImageNet test set. The power data for QKFormer and ANNs is evaluated according to Appendix.6.6, and the power data for other works were obtained from related papers. "A2S" denotes "ANN-to-SNN", "HST-$L$-$D$" denotes "Hierarchical Spiking Transformer" with $L$ encoder blocks and $D$ channels. HST-10-768 and HST-10-768∗∗ means HST-10-768 with 2882 and 3842 input size for inference. The top-5 accuracy of QKFormer (HST-10-768∗∗) is 97.74%.
@@
RMP[21] A2S VGG-16 2242 39.90 - 2048 73.09
QCFS[22] A2S ResNet-18 2242 11.70 - 1024 74.32
MST[23] A2S Swin Transformer-T 2242 28.50 - 512 78.51
SEW ResNet[28] SNN SEW-ResNet-34 2242 21.79 4.89 4 67.04
SNN SEW-ResNet-101 2242 44.55 8.91 4 68.76
SNN SEW-ResNet-152 2242 60.19 12.89 4 69.26
Spikformer[11] SNN Spikformer-8-384 2242 16.81 7.73 4 70.24
SNN Spikformer-8-512 2242 29.68 11.58 4 73.38
SNN Spikformer-8-768 2242 66.34 21.48 4 74.81
Spikingformer[12] SNN Spikingformer-8-384 2242 16.81 4.69 4 72.45
SNN Spikingformer-8-512 2242 29.68 7.46 4 74.79
SNN Spikingformer-8-768 2242 66.34 13.68 4 75.85
S-Transformer[13] SNN S-Transformer-8-384 2242 16.81 3.90 4 72.28
SNN S-Transformer-8-512 2242 29.68 1.13 1 71.68
SNN S-Transformer-8-512 2242 29.68 4.50 4 74.57
SNN S-Transformer-8-768 2882 66.34 6.09 4 77.07
ViT[4] ANN ViT-B/16 3842 86.59 254.84 1 77.90
DeiT[32] ANN DeiT-B 2242 86.59 80.50 1 81.80
ANN DeiT-B 3842 86.59 254.84 1 83.10
Swin[8] ANN Swin Transformer-B 2242 87.77 70.84 1 83.50
ANN Swin Transformer-B 3842 87.77 216.20 1 84.50
QKFormer SNN HST-10-384 2242 16.47 15.13 4 78.80
SNN HST-10-512 2242 29.08 21.99 4 82.04
SNN HST-10-768 2242 64.96 8.52 1 81.69
SNN HST-10-768 2242 64.96 38.91 4 84.22
SNN HST-10-768 2882 64.96 64.27 4 85.25
SNN HST-10-768∗∗ 3842 64.96 113.64 4 85.65
@@ 4.1 Results on ImageNet-1k Classification
Experimental Setup on ImageNet. In this experiment, we use AdamW as the optimizer, which is adopted with a base learning rate of $6\times 10^{-4}$. The actual learning rate was calculated as BatchSize/256 multiplied by the base learning rate. The batch size is set to $512$, which is realized by accumulated gradient iterations [33] and distributed across 8 Nvidia V100 GPUs. We trained QKFormer for $200$ epochs. In addition, following DeiT [32], data augmentation techniques including RandAugment [34], random erasing [35], and stochastic depth [36] are employed in this study. The number of blocks in the three stages is set as {1, 2, 7} respectively.
@@
Main Results on ImageNet. The experimental results demonstrate the superior performance of our proposed QKFormer, surpassing previous works’ performance by a large margin (Table 2). QKFormer (64.96 M) achieves 85.65% top-1 accuracy and 97.74% top-5 accuracy on ImageNet. To begin with, we compare our model with the baseline spiking transformer (i.e., Spikformer [11]). Our QKFormer models have slightly fewer parameters but much higher performance. For example, our QKFormer (64.96 M, 85.65%) significantly outperforms Spikformer (66.34 M, 74.81%) by 10.84%. In addition, compared with SDSA, our Q-K attention has lower computational complexity (shown in Table 1) and our QKFormer has much higher performance than S-Transformer (built by SDSA) [13]. In detail, QKFormer outperforms S-Transformer by 7.55%, 7.47%, and 8.58% respectively on three models with comparable #parameters. Finally, Our QKFormer outperforms the SOTA ANN-to-SNN model MST [23] by 7.14% and has much fewer time steps meanwhile. To our best knowledge, this is the first time that a direct training SNN model has achieved an accuracy of over 85% on ImageNet-1k.
Comparing with ANN Models on ImageNet. Our QKFormer is an event-driven SNN model, whose output is in binary form (either 0 or 1), the multiplications of activations and weights can be transformed into sparse addition, thus enjoying high energy efficiency. It should be noted that hierarchical architecture will lead to the power increment of QKFormer. This is still very cost-effective compared with ANN models. For instance, QKFormer (64.96M, 85.65%, SNN, 113.64mJ) Vs. Swin Transformer (88M, 84.5%, ANN, 216.20mJ) [8] Vs. DeiT-B (86M, 83.1%, ANN, 254.84mJ) [32] Vs. ViT (85.59M, 77.9%, ANN, 254.84mJ). [4]. Under the same experiment conditions without pre-training or extra training data, our QKFormer has surpassed the most well-known Transformer-based ANNs in performance while maintaining high energy efficiency.
Table 3: Comparision on CIFAR10, CIFAR100, DVS128, CIFAR10-DVS. "Param" denotes "Parameter (M)", "Acc" denotes "Top-1 Accuracy (%)", "$T$" denotes "Time Step".
Method CIFAR10 CIFAR100 DVS128 CIFAR10-DVS
Param $T$ Acc Param $T$ Acc Param $T$ Acc Param $T$ Acc
Spikformer [11] 9.32 4 95.51 9.32 4 78.21 2.57 16 98.3 2.57 16 80.9
Spikingformer [12] 9.32 4 95.81 9.32 4 78.21 2.57 16 98.3 2.57 16 81.3
CML [14] 9.32 4 96.04 9.32 4 80.02 2.57 16 98.6 2.57 16 80.9
S-Transformer[13] 10.28 4 95.60 10.28 4 78.4 2.57 16 99.3 2.57 16 80.0
STSA[15] -- -- -- -- -- -- 1.99 16 98.7 1.99 16 79.93
ResNet-19 (ANN) 12.63 1 94.97 12.63 1 75.35 -- -- -- -- -- --
Trasnformer (ANN) 9.32 1 96.73 9.32 1 81.02 -- -- -- -- -- --
QKFormer 6.74 4 96.18 6.74 4 81.15 1.50 16 98.6 1.50 16 84.0
4.2 Results on CIFAR and Neuromorphic Datasets
@@
CIFAR Classification. In this experiment, the QKFormer is trained for 400 epochs with a batch size of 64 following previous works: Spikformer [11], Spikingformer [12]. Following Spikformer, we use 4 blocks in QKFormer in total, which are distributed {1, 1, 2} in three stages. Due to the hierarchical architecture design, our QKFormer model has only 6.74 M parameters in this case. The results on CIFAR datasets are shown in Table 3. For CIFAR10, our model achieved 96.18% accuracy with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 0.67% and reduces 2.58 M parameters meanwhile. For CIFAR100, our model achieved 81.15% with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 2.94% and reduces 2.58 M parameters meanwhile.
@@
Neuromorphic Classification. We compare our method with SOTA methods on both CIFAR10-DVS and DVS-Gesture datasets. In this experiment, We utilize a mini QKFormer model with 1.50 M parameter, which has {0, 1, 1} blocks in three stages. The max patch embedding dimension is set to 256. The training process involves 200 epochs for DVS128 Gesture and 106 epochs for CIFAR10-DVS. The number of time steps of the spiking neuron is 10 or 16. The experimental results of temporal neuromorphic classification are presented in Table 3. For DVS128-Gesture dataset, our model with 1.50 M parameters achieves 98.6% accuracy using 16 time steps and 98.3% accuracy using 10 time steps. For CIFAR10-DVS dataset, our model achieves 84.0% accuracy with only 1.50 M parameters using 16 time steps. Our proposed QKFormer significantly outperforms Spikformer by 3.1% while reducing 1.07 M parameters. In addition, our model with 10 time steps achieves 83.8% accuracy, which outperforms Spikformer by 4.9% and outperforms the SOTA model (Spikingformer) by 3.9%.
Figure 3: The visualization and memory consumption of QKTA. 3 is the visualization of Q-K token attention. The white dot means value 1, while the black one means value 0. 3 shows the comparison of memory costs between QKTA and SSA under different token numbers. $N$ is the token number.
Figure 4: (a) shows the variance and expectation of SSA, (b) shows the variance and expectation of QKTA. Assume that all the spike elements (either 0 or 1) in SSA and QKTA are independent random variables and subject to Bernoulli distribution.